Two angles are supplementary, and one is 5 degrees more than six times the other. Find the larger supplementary angle?

Answer by checkley79(3341)

x+y=180
x=6y+5
6y+5+y=180
7y=180-5
7y=175
y=175/7
y=25 ans. for the smaller angle.
x=6*25+5=150+5=155 ans. for the larger angle.
Proof:
155+25=180
180=180